An aqueous solution has a mole fraction of 0.068 KOH. What is the molarity of the solution? Assume that the density of the solution is 1.15 g/mL.
Answer #1

Let the volume of solution = 1 L

Mass of solution = density×volume = 1.15×1000 = 1150 g

Let mass of KOH = x g

Mass of water = 1150-x

Molefraction of KOH = moles of KOH/total moles

0.068 = x/56/(x/56 + (1150-x)/18)

1 + (1150-x)56/18x = 1/0.068

1150×56/18x - 56/18 = 14.706

3577.77/x - 3.11 = 14.706

3577.77/x = 17.816

x = 3577.77/17.816

= 200.82 g

Molarity of KOH = mass/molar mass×volume

= 200.82/56×1

= 3.59 M